The formula that describes an object’s motion is given by S=ut+12at2
S
=
u
t
+
1
2
a
t
2
, where S
S
is the distance traveled, u
u
is the initial velocity, a
a
is the acceleration, and t
t
is the time. Which equation represents a
a
in terms of the other variables?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

s=ut+(1/2)at^2
2(s-ut)=at^2
a=2(s-ut)/t^2. This may be written in other ways.
the units are:
m=(m/sec)*sec + a* sec^2
a has to be m/sec^2
m/sec^2=(m-m/sec*sec)/sec^2=m/sec^2

Answer by ikleyn(52795)

.
Please

d
o

n
o
t

p
o
s
t

i
n

t
h
i
s

f
o
r
m
a
t